OpenStudy (anonymous):

MEDAL & FAN! AND SHOUTOUT IF YOU WANT IT! What is the length of a side of a cube with a volume of 2197 ft3?

OpenStudy (gorv):

take its cue root

OpenStudy (gorv):

volume of cube=a^3 a^3=2197 a=(2197)^1/3

OpenStudy (anonymous):

Thanks but I just need an answer... I f you could explain it in words for me to figure out that would be better...

OpenStudy (gorv):

a=13

OpenStudy (rock_mit182):